Philips Curve.
I = 36 - 4U.
I = 12.
Determine the unemployment rate.?
1
Expert's answer
2016-05-12T12:36:03-0400
If the equation of the Philips Curve is I = 36 - 4U and I = 12, then the unemployment rate U is:
12 = 36 - 4U,
4U = 24,
U = 6%.